ROTATIONAL SPECTROSCOPY OF THIOFORMALDEHYDE, H_{2}CS, IN ITS FOUR LOWEST EXCITED VIBRATIONAL, CORIOLIS-COUPLED STATES

\underline{HOLGER S. P. MÜLLER}, CHRISTIAN P. ENDRES, STEPHAN SCHLEMMER, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; ATSUKO MAEDA, IVAN MEDVEDEV, ERIC HERBST, MANFRED WINNEWISSER, FRANK C. De LUCIA, Department of Physics, The Ohio State University, Columbus, OH 43210, USA.

Rotational spectra of vibrationally excited states of H_{2}CS below 1500 cm^{−1} have been obtained by long-path absorption spectroscopy between 120 and currently 670 GHz. The highest K_{a} accessed range from 9 to 12 and correspond to combined rotational and vibrational energies of almost 2400 cm^{−1}. The v_{4} = 1 (990.185 cm^{−1}) and v_{6} = 1 (991.019 cm^{−1}) states are almost degenerate and undergo very strong Coriolis interaction (|ζ_{46}^{a}| = 0.5148) which causes K_{a} > 0 to be mixed completely. The v_{3} = 1 state is at 1059.204 cm^{−1} and the first order Coriolis terms are smaller (|ζ_{36}^{c}| = 0.2743 and |ζ_{34}^{b}| = 0.0574) so that perturbations are more local. Finally, the v_{2} = 1 state (1455.496 cm^{−1}) is comparatively far away from the three lower states, but there is a large coupling term with v_{4} = 1: |ζ_{24}^{b}| = 0.8555. Because of the large energy difference, only higher order effects are pronounced in this case. The initial assignments were greatly facilitated by previous infrared studies on ν_{4}, ν_{6}, and ν_{3}^{a}^{b} as well as on ν_{2}.^{c} The results from these as well as previous and present pure rotational studies on the ground vibrational state were considered in the global fit. The very satifactory fit required a modest number of vibrational corrections for the four vibrationally excited states and a comparatively large number of Coriolis interaction terms.

aD. J. Bedwell and G. Duxbury, J. Mol. Spectrosc. 84 (1980) 531−558.

bP. H. Turner, L. Halonen, and I. M. Mills, J. Mol. Spectrosc. 88 (1981) 402−419.

cD. McNaughton and D. N. Bruget, J. Mol. Spectrosc. 159 (1993) 340−349.